# Generalization Below the Edge of Stability: The Role of Data Geometry

### v1

@@ F.2.1 Analysis of the Global Weight Function
For any mixed distribution $\mathcal{P}_{{\bm{X}}}=\sum_{j=1}^{J}p_{j}\mathcal{P}_{{\bm{X}},j}$ with $\mathrm{supp}(\mathcal{P}_{{\bm{X}},j})=V_{j}$. Let $g$ be the global weight induced by the mixture $\mathcal{P}_{{\bm{X}}}$, and $g_{j}$ the weight induced by $\mathcal{P}_{{\bm{X}},j}$. For every $j\in\{1,\dots,J\}$,
@@
$\mathcal{F}_{g}(\mathbb{B}^{V_{j}}_{1};M,C)\;\subseteq\;\mathcal{F}_{g_{j}}\!\big(\mathbb{B}^{V_{j}}_{1};M,\sqrt{2}\,C/p_{j}^{2}\big)\,.$ (37)
@@
$g({\bm{u}},t)$ $=\mathcal{P}_{{\bm{X}}}(A)^{2}\cdot\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}}}}[{\bm{X}}^{\mathsf{T}}{\bm{u}}-t\mid A]\cdot\sqrt{1+\|\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}}}}[{\bm{X}}\mid A]\|_{2}^{2}}\,$
$g_{j}({\bm{u}},t)$ $=\mathcal{P}_{{\bm{X}}}(A\mid{\bm{x}}\in V_{j})^{2}\cdot\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}}}}[{\bm{X}}^{\mathsf{T}}{\bm{u}}-t\mid A,{\bm{x}}\in V_{j}]\cdot\sqrt{1+\|\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}}}}[{\bm{X}}\mid A,{\bm{x}}\in V_{j}]\|_{2}^{2}}\,$
@@
Using the law of total probability and total expectation for the mixture distribution $\mathcal{P}_{{\bm{X}}}=\sum_{i=1}^{J}p_{i}\mathcal{P}_{{\bm{X}},i}$, and the non-negativity of $({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}$, we get
$\mathcal{P}_{{\bm{X}}}(A)\;\geq\;p_{j}\,\mathcal{P}_{{\bm{X}},j}(A),\qquad\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}}}}[({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}]\;\geq\;p_{j}\,\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}},j}}[({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}].$
Hence, by combining the first two terms of $g({\bm{u}},t)$ as $\mathcal{P}_{{\bm{X}}}(A)\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}}}}[({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}]$, we have:
@@
Let $\mathcal{P}$ be a distribution defined in Assumption 3.1 and recall that $\mathcal{P}_{j}$ is $\mathcal{P}$ conditional to ${\bm{x}}\in V_{j}$. Fix $j\in\{1,\dots,J\}$ and a data set $\mathcal{D}\sim\mathcal{P}^{\otimes n}$. Let $\mathcal{D}_{j}:=\mathcal{D}\cap V_{j}$ and $n_{j}:=|\mathcal{D}_{j}|$. Then with probability $1-\delta$,
@@ Appendix G Generalization Upper Bounds: Isotropic Beta Family
Let $\bm{X}$} be a $d$-dimensional random vector in $\mathbb{R}^{d}$. For any $\alpha\in(0,\infty)$, the isotropic $\alpha$-powered-radial distribution is defined by the generation process
${\bm{X}}=h(R){\bm{U}}\sim\mathcal{P}_{{\bm{X}}}(\alpha),$ (43)
@@
Let $\mathcal{P}_{{\bm{X}}}(\alpha)$ be the isotropic $\alpha$-powered-radial distribution in Definition G.1. For ${\bm{X}}\sim\mathcal{P}_{{\bm{X}}}(\alpha)$ any $t\in[0,1]$, $\mathbb{P}\left(\|{\bm{X}}\|>1-t\right)=t^{\alpha}$. In particular, $\|{\bm{X}}\|_{2}$ is a $\mathrm{Beta}(1,\alpha)$ distribution.
@@ G.1 Characterization of the Weight Function for a Custom Radial Distribution
In this section, we analyze the properties of the weight function $g_{\alpha}({\bm{u}},t)=g_{\mathcal{P}_{\vec{X},\alpha}}({\bm{u}},t)$ with respect to the population distribution $\mathcal{P}_{\vec{X},\alpha}$ we defined in Definition G.1 and Assumption G.3. Recall that $g_{\alpha}({\bm{u}},t)=\min\left(\tilde{g}_{\alpha}({\bm{u}},t),\tilde{g}_{\alpha}(-{\bm{u}},-t)\right)$, where
@@
For any fixed $\varepsilon<1$, we may decompose $\mathbb{B}_{1}^{d}$ into $\varepsilon$-annulus $\mathbb{A}^{d}_{\varepsilon}\coloneqq\{{\bm{x}}\in\mathbb{B}_{1}^{d}\mid\|{\bm{x}}\|_{2}\geq 1-\varepsilon\}$ and the closure of its complement is called $\varepsilon$-strict interiordenoted by $\mathds{1}^{d}_{\varepsilon}$.
@@
Using the property that the marginal distribution of $\bm{x}$} is $\mathcal{P}_{{\bm{X}}}(\alpha)$ and its concentration property, with probability at least $1-\delta$,
@@ H.1 Construction of “hard-to-learn” networks
The $L^{2}(\mathcal{P}_{\bm{X}}(\alpha))$-norm of $\varphi_{{\bm{u}},\varepsilon^{2}}$, where the measure $\mathcal{P}_{\bm{X}}(\alpha)$ is defined in Definition G.1, is given by
@@
    • Radial Measure: The condition requires the radius $r$ to be near $1$. For the $\mathcal{P}_{{\bm{X}}}(\alpha)$ distribution, this confines $r$ to a region of length $\Delta r\sim O(\varepsilon^{2\alpha})$.
@@
The squared $L^{2}$ norm of $\varphi_{{\bm{u}},\varepsilon^{2}}$ over the distribution $\mathcal{P}_{{\bm{X}}}(\alpha)$ is defined by the expectation
$I=\|\varphi_{{\bm{u}},\varepsilon^{2}}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}=\mathbb{E}_{{\bm{X}}\sim\mathcal{P}_{{\bm{X}}}(\alpha)}\left[|\varphi_{{\bm{u}},\varepsilon^{2}}({\bm{X}})|^{2}\right]$
@@
There exist constants depending only on $(d,\alpha)$, such that $\mathcal{P}_{{\bm{X}}}\big(C({\bm{u}},\varepsilon)\big)\ \asymp\varepsilon^{\,d-1+2\alpha}$.
@@
As the atoms $\Phi_{i}$ have disjoint supports, the squared $L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))$ distance between any two distinct functions $f_{\xi}$ and $f_{\xi^{\prime}}$ can be computed as:
$\|f_{\xi}-f_{\xi^{\prime}}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}=\sum_{i=1}^{N}(\xi_{i}-\xi^{\prime}_{i})^{2}\|\Phi_{i}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}=4\sum_{i:\xi_{i}\neq\xi^{\prime}_{i}}\|\Phi_{i}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}.$
Referring to the cap mass properties in Lemma H.1 (which implies $\|\Phi_{i}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\asymp\varepsilon^{d-1+2\alpha}$), this simplifies to the final distance scaling
$\|f_{\xi}-f_{\xi^{\prime}}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\ \asymp_{d,\alpha}\ \varepsilon^{\,d-1+2\alpha}\,d_{H}(\xi,\xi^{\prime}),$
@@
Let $\{C({\bm{u}}_{i},\varepsilon)\}_{i=1}^{N}$ be the set of disjoint caps from Lemma H.3. Let the sample size be $N_{\text{poi}}\sim\mathrm{Poi}(n)$. Let $Z_{i}$ be the number of samples falling into cap $C({\bm{u}}_{i},\varepsilon)$. Define the expected number of points per cap as $\lambda:=n\cdot\mathcal{P}_{{\bm{X}}}(C({\bm{u}}_{1},\varepsilon))$. If we choose $\varepsilon$ such that $\lambda\asymp 1$, then there exists a constant $c>0$ such that with probability at least $1-\exp(-cN)$:
@@
Under Poissonization, the random variables $Z_{i}=\#\{{\bm{x}}_{j}\in C({\bm{u}}_{i},\varepsilon)\}$ are independent Poisson variables with mean $\lambda_{i}=n\cdot\mathcal{P}_{{\bm{X}}}(C({\bm{u}}_{i},\varepsilon))$. By Lemma H.2 and our choice of scale, $\lambda_{i}=\lambda\asymp 1$ for all $i$.
@@
    Separation in Population: $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\asymp n^{-\frac{2\alpha}{\,d-1+2\alpha\,}}$.
@@
    $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\asymp\varepsilon^{\,d-1+2\alpha}\cdot d_{H}(\xi,\xi^{\prime})\asymp\varepsilon^{\,d-1+2\alpha}\cdot N\asymp\varepsilon^{\,d-1+2\alpha}\cdot\varepsilon^{-(d-1)}=\varepsilon^{2\alpha}.$
@@
    $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\asymp\left(n^{-1/(d-1+2\alpha)}\right)^{2\alpha}=n^{-\frac{2\alpha}{\,d-1+2\alpha\,}}.$
@@
$\inf_{\hat{f}}\sup_{f_{0}\in\mathcal{F}_{\text{pack}}}\mathbb{E}\left[\|\hat{f}-f_{0}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\right]\gtrsim n^{-\frac{2\alpha}{d-1+2\alpha}}.$
@@
Since on the event $E$, the separation $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\asymp n^{-\frac{2\alpha}{d-1+2\alpha}}$ and $\mathbb{P}(E)\to 1$ as $n\to\infty$, the result follows. ∎
@@
Let $\mathcal{P}$ denote any joint distribution of $({\bm{x}},y)$ where the marginal distribution of $\bm{x}$} is $\mathcal{P}_{{\bm{X}}}(\alpha))$ and $y$ is supported on [1,1][-1,1]. Let $\mathcal{D}_{n}=\{({\bm{x}}_{j},y_{j})\}_{j=1}^{n}$ be a dataset of $n$ i.i.d. samples from $\mathcal{P}$. Let $\widehat{R}_{\mathcal{D}_{n}}(f)$ be any empirical risk estimator for the true risk $R_{\mathcal{P}}(f):=\mathbb{E}_{({\bm{x}},y)\sim\mathcal{P}}[(f({\bm{x}})-y)^{2}]$. Then,
@@
$=\frac{1}{2}\inf_{\hat{f}}\sup_{f_{0}\in\mathcal{F}_{\text{pack}}}\mathbb{E}\left[\|\hat{f}-f_{0}\|_{L^{2}(\mathcal{P}_{{\bm{X}}}(\alpha))}^{2}\right]$ (83)

### v2

@@ F.2.1 Analysis of the Global Weight Function
For any mixed distribution $\mathcal{P}_{X}=\sum_{j=1}^{J}p_{j}\mathcal{P}_{{\bm{X}},j}$ with $\mathrm{supp}(\mathcal{P}_{{\bm{X}},j})=V_{j}$. Let $g$ be the global weight induced by the mixture $\mathcal{P}_{X}$, and $g_{j}$ the weight induced by $\mathcal{P}_{{\bm{X}},j}$. For every $j\in\{1,\dots,J\}$,
@@
$\mathcal{F}_{g}(\mathbb{B}^{V_{j}}_{1};M,C)\;\subseteq\;\mathcal{F}_{g_{j}}\!\big(\mathbb{B}^{V_{j}}_{1};M,\sqrt{2}\,C/2p_{j}^{2}\big)\,.$ (41)
@@
$g({\bm{u}},t)$ $=\mathcal{P}_{X}(A)^{2}\cdot\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{X}}[{\bm{X}}^{\mathsf{T}}{\bm{u}}-t\mid A]\cdot\sqrt{1+\|\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{X}}[{\bm{X}}\mid A]\|_{2}^{2}}\,$
$g_{j}({\bm{u}},t)$ $=\mathcal{P}_{X}(A\mid{\bm{x}}\in V_{j})^{2}\cdot\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{X}}[{\bm{X}}^{\mathsf{T}}{\bm{u}}-t\mid A,{\bm{x}}\in V_{j}]\cdot\sqrt{1+\|\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{X}}[{\bm{X}}\mid A,{\bm{x}}\in V_{j}]\|_{2}^{2}}\,$
@@
Using the law of total probability and total expectation for the mixture distribution $\mathcal{P}_{X}=\sum_{i=1}^{J}p_{i}\mathcal{P}_{{\bm{X}},i}$, and the non-negativity of $({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}$, we get
$\mathcal{P}_{X}(A)\;\geq\;p_{j}\,\mathcal{P}_{{\bm{X}},j}(A),\qquad\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{X}}[({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}]\;\geq\;p_{j}\,\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{{\bm{X}},j}}[({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}].$
Hence, by combining the first two terms of $g({\bm{u}},t)$ as $\mathcal{P}_{X}(A)\mathop{\mathbb{E}}_{{\bm{x}}\sim\mathcal{P}_{X}}[({\bm{X}}^{\mathsf{T}}{\bm{u}}-t)\mathds{1}_{A}]$, we have:
@@
Let $\mathcal{P}$ be a distribution defined in Assumption F.1 and recall that $\mathcal{P}_{j}$ is $\mathcal{P}$ conditional to ${\bm{x}}\in V_{j}$. Fix $j\in\{1,\dots,J\}$ and a data set $\mathcal{D}\sim\mathcal{P}^{\otimes n}$. Let $\mathcal{D}_{j}:=\mathcal{D}\cap V_{j}$ and $n_{j}:=|\mathcal{D}_{j}|$. Then with probability $1-\delta$,
@@ Appendix G Generalization Upper Bounds: Isotropic Beta Family
Let $\bm{X}$} be a $d$-dimensional random vector in $\mathbb{R}^{d}$. For any $\alpha\in(0,\infty)$, the isotropic $\mathrm{Beta}(\alpha)$-radial distribution is defined by the generation process
${\bm{X}}=h(R){\bm{U}}\sim\mathcal{P}_{X}(\alpha),$ (47)
@@
Let $\mathcal{P}_{X}(\alpha)$ be the isotropic $\mathrm{Beta}(\alpha)$-radial distribution in Definition 3.2. For ${\bm{X}}\sim\mathcal{P}_{X}(\alpha)$ any $t\in[0,1]$, $\mathbb{P}\left(\|{\bm{X}}\|>1-t\right)=t^{\alpha}$. In particular, $\|{\bm{X}}\|_{2}$ is a $\mathrm{Beta}(1,\alpha)$ distribution.
@@ G.1 Characterization of the Weight Function for a Custom Radial Distribution
In this section, we analyze the properties of the weight function $g_{\alpha}({\bm{u}},t)=g_{\mathcal{P}_{\vec{X},\alpha}}({\bm{u}},t)$ with respect to the population distribution $\mathcal{P}_{\vec{X},\alpha}$ we defined in Definition 3.2 and Assumption G.3. Recall that $g_{\alpha}({\bm{u}},t)=\min\left(\tilde{g}_{\alpha}({\bm{u}},t),\tilde{g}_{\alpha}(-{\bm{u}},-t)\right)$, where
@@
For any fixed $\varepsilon<1$, we may decompose $\mathbb{B}_{1}^{d}$ into $\varepsilon$-annulus $\mathbb{A}^{d}_{\varepsilon}\coloneqq\{{\bm{x}}\in\mathbb{B}_{1}^{d}\mid\|{\bm{x}}\|_{2}\geq 1-\varepsilon\}$ and the closure of its complement is called $\varepsilon$-strict interior denoted by $\mathbb{I}^{d}_{\varepsilon}=\mathbb{B}_{1-\varepsilon}^{d}$.
@@
Using the property that the marginal distribution of $\bm{x}$} is $\mathcal{P}_{X}(\alpha)$ and its concentration property, with probability at least $1-\delta$,
@@ H.1 Construction of “hard-to-learn” networks
The $L^{2}(\mathcal{P}_{\bm{X}}(\alpha))$-norm of $\varphi_{{\bm{u}},\varepsilon^{2}}$, where the measure $\mathcal{P}_{\bm{X}}(\alpha)$ is defined in Definition 3.2, is given by
@@
    • Radial Measure: The condition requires the radius $r$ to be near $1$. For the $\mathcal{P}_{X}(\alpha)$ distribution, this confines $r$ to a region of length $\Delta r\sim O(\varepsilon^{2\alpha})$.
@@
The squared $L^{2}$ norm of $\varphi_{{\bm{u}},\varepsilon^{2}}$ over the distribution $\mathcal{P}_{X}(\alpha)$ is defined by the expectation
$I=\|\varphi_{{\bm{u}},\varepsilon^{2}}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}=\mathbb{E}_{{\bm{X}}\sim\mathcal{P}_{X}(\alpha)}\left[|\varphi_{{\bm{u}},\varepsilon^{2}}({\bm{X}})|^{2}\right]$
@@
There exist constants depending only on $(d,\alpha)$, such that $\mathcal{P}_{X}\big(C({\bm{u}},\varepsilon)\big)\ \asymp\varepsilon^{\,d-1+2\alpha}$.
@@
As the atoms $\Phi_{i}$ have disjoint supports, the squared $L^{2}(\mathcal{P}_{X}(\alpha))$ distance between any two distinct functions $f_{\xi}$ and $f_{\xi^{\prime}}$ can be computed as:
$\|f_{\xi}-f_{\xi^{\prime}}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}=\sum_{i=1}^{N}(\xi_{i}-\xi^{\prime}_{i})^{2}\|\Phi_{i}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}=4\sum_{i:\xi_{i}\neq\xi^{\prime}_{i}}\|\Phi_{i}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}.$
Referring to the cap mass properties in Lemma H.1 (which implies $\|\Phi_{i}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\asymp\varepsilon^{d-1+2\alpha}$), this simplifies to the final distance scaling
$\|f_{\xi}-f_{\xi^{\prime}}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\ \asymp_{d,\alpha}\ \varepsilon^{\,d-1+2\alpha}\,d_{H}(\xi,\xi^{\prime}),$
@@
Let $\{C({\bm{u}}_{i},\varepsilon)\}_{i=1}^{N}$ be the set of disjoint caps from Lemma H.3. Let the sample size be $N_{\text{poi}}\sim\mathrm{Poi}(n)$. Let $Z_{i}$ be the number of samples falling into cap $C({\bm{u}}_{i},\varepsilon)$. Define the expected number of points per cap as $\lambda:=n\cdot\mathcal{P}_{X}(C({\bm{u}}_{1},\varepsilon))$. If we choose $\varepsilon$ such that $\lambda\asymp 1$, then there exists a constant $c>0$ such that with probability at least $1-\exp(-cN)$:
@@
Under Poissonization, the random variables $Z_{i}=\#\{{\bm{x}}_{j}\in C({\bm{u}}_{i},\varepsilon)\}$ are independent Poisson variables with mean $\lambda_{i}=n\cdot\mathcal{P}_{X}(C({\bm{u}}_{i},\varepsilon))$. By Lemma H.2 and our choice of scale, $\lambda_{i}=\lambda\asymp 1$ for all $i$.
@@
    Separation in Population: $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\asymp n^{-\frac{2\alpha}{\,d-1+2\alpha\,}}$.
@@
    $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\asymp\varepsilon^{\,d-1+2\alpha}\cdot d_{H}(\xi,\xi^{\prime})\asymp\varepsilon^{\,d-1+2\alpha}\cdot N\asymp\varepsilon^{\,d-1+2\alpha}\cdot\varepsilon^{-(d-1)}=\varepsilon^{2\alpha}.$
@@
    $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\asymp\left(n^{-1/(d-1+2\alpha)}\right)^{2\alpha}=n^{-\frac{2\alpha}{\,d-1+2\alpha\,}}.$
@@
$\inf_{\hat{f}}\sup_{f_{0}\in\mathcal{F}_{\text{pack}}}\mathbb{E}\left[\|\hat{f}-f_{0}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\right]\gtrsim n^{-\frac{2\alpha}{d-1+2\alpha}}.$
@@
Since on the event $E$, the separation $\|f-f^{\prime}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\asymp n^{-\frac{2\alpha}{d-1+2\alpha}}$ and $\mathbb{P}(E)\to 1$ as $n\to\infty$, the result follows. ∎
@@
Let $\mathcal{P}$ denote any joint distribution of $({\bm{x}},y)$ where the marginal distribution of $\bm{x}$} is $\mathcal{P}_{X}(\alpha))$ and $y$ is supported on [1,1][-1,1]. Let $\mathcal{D}_{n}=\{({\bm{x}}_{j},y_{j})\}_{j=1}^{n}$ be a dataset of $n$ i.i.d. samples from $\mathcal{P}$. Let $\widehat{R}_{\mathcal{D}_{n}}(f)$ be any empirical risk estimator for the true risk $R_{\mathcal{P}}(f):=\mathbb{E}_{({\bm{x}},y)\sim\mathcal{P}}[(f({\bm{x}})-y)^{2}]$. Then,
@@
$=\frac{1}{2}\inf_{\hat{f}}\sup_{f_{0}\in\mathcal{F}_{\text{pack}}}\mathbb{E}\left[\|\hat{f}-f_{0}\|_{L^{2}(\mathcal{P}_{X}(\alpha))}^{2}\right]$ (87)